# Tropical curves of unibranch points and hypertangency

### v2

@@ 4.1. Hypertangency at unibranch point
[…]
@@
by the induction hypothesis ($h-2$ is even). The case $h$ odd follows in the same way. The Lemma is proved. ∎
@@ 4.2. Intersection multiplicity at equivalent points
If $n/m\in\mathbb{N}$ there is nothing left to prove. Let $n/m\not\in\mathbb{N}$, hence $k<n/m$. Then $C^{k}$ has an $(n-m,m)$-point w.r.t. $E_{k}$, and $B^{k}$ has an $(n^{\prime}-m^{\prime},m^{\prime})$-point w.r.t. $E_{k}$ (by Lemma 2.1.2); let us show that this point is $q^{k}$ for both $C^{k}$ and $B^{k}$. We have
@@ 5.1. Genus computation
where $p$ varies over all points infinitely near to $q$, and $m_{p}$ is the multiplicity of $p$. For example, if $q$ is an ordinary $m$-fold point (i.e. $C$ has a $m$ branches meeting transversally in $q$), then $\delta(q)=m(m-1)/2$.
@@
In our case $l_{0}=n$,   $l_{1}=m$$l_{r}=c$ and $l_{r+1}=0$, from which (20) follows. To prove (21) we proceed by induction on $r$. If $r=1$ then
@@ 5.2. Dimension computation
(23) $m(m+1)/2-m=m(m-1)/2.$
Next, $C$ is unibranch at $q$, and its tangent line there is $y=0$. Therefore the only summand of degree $m$ is $a_{0,m}y^{m}$ and we get the following new $m-1$ conditions:
@@
with $j=l$ and $i=h-l+m$. The conditions imposed earlier imply exactly that $f^{\prime}$ has no summand of type $x^{h}$ for all $h=0,\ldots,d-m-1$, and no summand of type $v^{l}$ for all $l=1,\ldots,m-1$.
@@
Let $n>2m$. Then $m=m^{\prime}$ and $C^{\prime}$ has a an $(m,n-m)$-point at $q^{\prime}$ with respect to $L^{\prime}$, by Lemma 2.1.2. Hence in $f^{\prime}$ there are no monomials of degree less than $m$, and no monomials of degree $m$, except $v^{m}$. These are $m(m+1)/2+m$ conditions, some of which are already satisfied, indeed $f^{\prime}$ has no term of type $x^{l}$ for all $l=0,\ldots,m$ ($m+1$ conditions), and no summand of type $v^{n}$ for all $n=1,\ldots,m-1$ ($m-1$ conditions). Hence the number of new conditions is
(26) $m(m+1)/2+m-(m+1)-(m-1)=m(m-1)/2.$
Let $n=2m$. Then $C^{\prime}$ has a an $m$-fold point at $q^{\prime}$. Hence in $f^{\prime}$ there are no monomials of degree less than $m$. Moreover, as we saw in the proof of Lemma 2.1.2, the homogeneous part of degree $m$ of $f^{\prime}$ has the form $(\alpha x+\beta v)^{m}$ with $\alpha^{m}=a_{2m,0}$ and $\beta^{m}=a_{0,m}$. This imposes $m-1$ conditions. We get a total of $m(m+1)/2+m-1$ conditions, but, as in the earlier cases, some of them already hold. In fact $f^{\prime}$ has no term of type $x^{l}$ for all $l=0,\ldots,m-1$ ($m$ conditions), and no summand of type $v^{n}$ for all $n=1,\ldots,m-1$ ($m-1$ conditions). Hence the number of new conditions is
(27) $m(m+1)/2+m-1-m-(m-1)=m(m-1)/2.$
@@
Now we need to apply the claim to all the curves appearing in the chain described in Proposition 2.1.4, of which we use the notation. Proposition 2.1.4 gives a sequence of curves whose multiplicity at the point lying over $q$ is equal to $l_{j}$, with $j=1,\ldots,r$. By the claim, each of these gives $l_{j}(l_{j}-1)/2$ new conditions. If $j\geq 2$ we get exactly $k_{j}$ such curves, hence a total of $k_{j}l_{j}(l_{j}-1)/2$ conditions. For $j=1$ we have only $k_{1}-1$ such blow-ups, hence $(k_{1}-1)m(m-1)/2$ conditions. But in the first part of the proof we had $m(m-1)/2$ conditions; see (23). Hence also for $j=1$ we have $k_{1}m(m-1)/2$ conditions, hence

### v3

@@ 4.1. Hypertangency at unibranch point
We shall prove $(m^{\prime},n^{\prime})\sim(m,n)$. We can assume $r\leq r^{\prime}$. Consider the chain of blow-ups described before Proposition 2.1.4 for the curve $C$. For every $j=1,\ldots,r$ and $i=1,\ldots,k_{j}$ we have a sequence of blow-ups, $S^{i}_{j}$, of $\mathbb{P}^{2}$, containing the proper transform, $C^{i}_{j}$, of $C$, and a point $q^{i}_{j}\in C^{i}_{j}$, which is the center of the blow up of $S^{i}_{j}$; we have $q=q^{1}_{1}$, $C=C^{1}_{1}$ and $\mathbb{P}^{2}=S^{1}_{1}$. We consider also the curve $C_{*}$ in the blow-up $S_{*}$ of $S^{k_{r}}_{r}$ at $q^{k_{r}}_{r}$, defined before Lemma 2.2.2.
@@
by the induction hypothesis (h2h-2 is even). The case $h$ odd follows in the same way. The Lemma is proved. ∎
@@
Example 4.1.4.
Let $B\subset\mathbb{P}^{2}$ be the projective conic of affine equation $y=x^{2}$, and $C\subset\mathbb{P}^{2}$ be the projective cubic of affine equation $y=x^{2}+y^{3}$. A simple computation shows that $B$ and $C$ intersect only at $q=(0,0)$. The theorem implies that $q$ is not a flex of $C$ (which is easy to check in this case).
@@ 4.2. Intersection multiplicity at equivalent points
If $n/m\in\mathbb{N}$ there is nothing left to prove. Let $n/m\not\in\mathbb{N}$, hence $k<n/m$. Then $C^{k}$ has an (nm,m)(n-m,m)-point w.r.t. $E_{k}$, and $B^{k}$ has an $(n^{\prime}-m^{\prime},m^{\prime})$-point w.r.t. $E_{k}$ (by Lemma 2.1.2); let us show that this point is $q^{k}$ for both $C^{k}$ and $B^{k}$. We have
@@ 5.1. Genus computation
where $p$ varies over all points infinitely near to $q$ (the points added over $q$ in the proper transforms of the blow-ups), and $m_{p}$ is the multiplicity of $p$. For example, if $q$ is an ordinary $m$-fold point (i.e. $C$ has $m$ branches meeting transversally in $q$), then $\delta(q)=m(m-1)/2$.
@@
In our case $l_{0}=n$, $l_{1}=m$, $l_{r}=c$ and $l_{r+1}=0$, from which (20) follows. To prove (21) we proceed by induction on $r$. If $r=1$ then
@@ 5.2. Dimension computation
(23) m(m+1)/2m=m(m1)/2.m(m+1)/2-m=m(m-1)/2.
Next, $C$ is unibranch at $q$, and its tangent line there is $y=0$. Therefore the only summand of degree $m$ is $a_{0,m}y^{m}$ and we get the following new m1m-1 conditions:
@@
with $j=l$ and i=hl+mi=h-l+m. The conditions imposed earlier imply exactly that $f^{\prime}$ has no summand of type $x^{h}$ for all $h=0,\ldots,d-m-1$, and no summand of type $v^{l}$ for all $l=1,\ldots,m-1$.
@@
Let $n>2m$. Then $m=m^{\prime}$ and $C^{\prime}$ has a an (m,nm)(m,n-m)-point at $q^{\prime}$ with respect to $L^{\prime}$, by Lemma 2.1.2. Hence in $f^{\prime}$ there are no monomials of degree less than $m$, and no monomials of degree $m$, except $v^{m}$. These are $m(m+1)/2+m$ conditions, some of which are already satisfied, indeed $f^{\prime}$ has no term of type $x^{l}$ for all $l=0,\ldots,m$ ($m+1$ conditions), and no summand of type $v^{n}$ for all $n=1,\ldots,m-1$ (m1m-1 conditions). Hence the number of new conditions is
(26) m(m+1)/2+m(m+1)(m1)=m(m1)/2.m(m+1)/2+m-(m+1)-(m-1)=m(m-1)/2.
Let $n=2m$. Then $C^{\prime}$ has a an $m$-fold point at $q^{\prime}$. Hence in $f^{\prime}$ there are no monomials of degree less than $m$. Moreover, as we saw in the proof of Lemma 2.1.2, the homogeneous part of degree $m$ of $f^{\prime}$ has the form $(\alpha x+\beta v)^{m}$ with $\alpha^{m}=a_{2m,0}$ and $\beta^{m}=a_{0,m}$. This imposes m1m-1 conditions. We get a total of m(m+1)/2+m1m(m+1)/2+m-1 conditions, but, as in the earlier cases, some of them already hold. In fact $f^{\prime}$ has no term of type $x^{l}$ for all $l=0,\ldots,m-1$ ($m$ conditions), and no summand of type $v^{n}$ for all $n=1,\ldots,m-1$ (m1m-1 conditions). Hence the number of new conditions is
(27) m(m+1)/2+m1m(m1)=m(m1)/2.m(m+1)/2+m-1-m-(m-1)=m(m-1)/2.
@@
Now we need to apply the claim to all the curves appearing in the chain described in Proposition 2.1.4, of which we use the notation. Proposition 2.1.4 gives a sequence of curves whose multiplicity at the point lying over $q$ is equal to $l_{j}$, with $j=1,\ldots,r$. By the claim, each of these gives $l_{j}(l_{j}-1)/2$ new conditions. If $j\geq 2$ we get exactly $k_{j}$ such curves, hence a total of $k_{j}l_{j}(l_{j}-1)/2$ conditions. For $j=1$ we have only $k_{1}-1$ such blow-ups, hence $(k_{1}-1)m(m-1)/2$ conditions. But in the first part of the proof we had m(m1)/2m(m-1)/2 conditions; see (23). Hence also for $j=1$ we have $k_{1}m(m-1)/2$ conditions, hence